Eye is most sensitive to 5500Å and the diameter of the pupil is about 2 mm. Find the
angular limit of resolution of eye.
1
Expert's answer
2019-08-26T10:42:23-0400

We can find the angular limit of resolution of eye from the formula for the diffraction limit to resolution (also known as Rayleigh criterion):


"\theta = 1.22 rad \cdot \dfrac{\lambda}{D},"

here, "\lambda" is the wavalength of the light, "D" is the diameter of the pupil.

Then, we get:


"\theta = 1.22 rad \cdot \dfrac{5500 \text{\AA} \cdot \dfrac{1m}{10^{10} \text{\AA}}}{2 mm \cdot \dfrac{1m}{10^3 mm}} = 3.35 \cdot 10^{-4} rad."

Answer:

"\theta = 3.35 \cdot 10^{-4} rad."